# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v2

@@ Appendix H Proof of Theorem 3
Our proof is similar to the proof of the lower bound for adversarial linear mixture MDPs with the unknown transition in Zhao et al. [2023]. We prove this lemma by reducing the MNL mixture MDP problem to a sequence of binary logistic bandit problems.
Specifically, we use each three layers to construct a block. Note that the third layer of block $i$ is also the first layer of block $i+1$ and hence there are total $H/2$ blocks. In each block, both the first and third layers of this block only have one state, and the second layer has two states. Here we take block $i$ as an example. The first two layers of this block are associated with transition probability $\mathbb{P}_{i,1}$ and $\mathbb{P}_{i,2}$. Denote by $s_{i,1}$ the only state in the first layer of this block. In the second layer of the block $i$, we assume there exist two states $s_{i,2}^{*}$ and $s_{i,2}$. Let $s_{i,3}$ be the only state in the third layer of this block. Further, for any $a\in\mathcal{A}$, let $\phi(s_{i,1},a,s_{i,2})=0$. The probability of transferring to state $s_{i,2}^{*}$ when executing action $a$ at state $s_{i,1}$ is $\rho_{a}$, in particular,
$\mathbb{P}_{i,1}(s_{i,2}^{*}\mid s_{i,1},a)$ $=\frac{\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a))^{\top}\theta_{i,1}^{*})}=\rho_{a},$
$\quad\mathbb{P}_{i,1}(s_{i,2}\mid s_{i,1},a)$ $=\frac{1}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}=1-\rho_{a}.$
For the second layer, the MDP instance satisfies that $\forall s=s_{i,2}^{*},s_{i,2}$, and $a\in\mathcal{A}$, $\mathbb{P}_{i,2}(s_{i,3}\mid s,a)=1$. The reward function satisfies $r_{k}(s_{i,1},a)=0$ for the first layer and $r_{k}(s_{i,2}^{*},a)=1$, $r_{k}(s_{i,2},a)=0$ for the second layer for all $a\in\mathcal{A}$.
Consider the logistic bandit problem where a learner selects action $x\in\mathbb{R}^{d}$ and receives a reward $r_{k}$ sampled from Bernoulli distributed with mean
$\mu(x^{\top}\theta^{*})=\frac{1}{1+\exp(x^{\top}\theta^{*})}.$
By this configuration, we can see that learning in this block of MDP can be regarded as learning a $d$-dimensional logistic bandit problem with $A$ arms, where the arm set is $\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$ and the expected reward of each arm is $\rho_{a}$. It is also clear that the optimal policy at state $s_{i,1}$ is to select action $a_{i,1}^{*}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$. Thus, learning this MNL mixture MDP equals to learning $H/2$ logistic bandit problems. This finishes the proof. ∎
Appendix I Supporting Lemmas
@@
$V_{t}=V+\sum_{s=1}^{t-1}X_{s}X_{s}^{\top}\quad S_{t}=\sum_{s=1}^{t-1}\eta_{s}X_{s}.$
Then, for any $\delta>0$, with probability at least $1-\delta$, for all $t\geq 0$,
$\left\|S_{t}\right\|_{V_{t}^{-1}}^{2}\leq 2R^{2}\log\left(\frac{\operatorname{det}\left(V_{t}\right)^{1/2}\operatorname{det}(V)^{-1/2}}{\delta}\right).$
@@
Suppose $x_{1},\ldots,x_{t}\in\mathbb{R}^{d}$ and for any $1\leq s\leq t$, $\lVert x_{s}\rVert_{2}\leq L$. Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$. Then, for any $1\leq s\leq t$, we have
$\det(V_{t})\leq\left(\lambda+\frac{tL^{2}}{d}\right)^{d}.$
@@
Let $Z$ be a positive define matrix and $\mathcal{X}$ be a convex set,
$\widetilde{\mathbf{x}}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\langle\mathbf{g},\mathbf{x}\rangle+\frac{1}{2\eta}\lVert\mathbf{x}-\widehat{\mathbf{x}}\rVert_{Z}^{2}.$
Then, for all $\mathbf{x}^{*}\in\mathcal{X}$, we have
$\lVert\widetilde{\mathbf{x}}-\widehat{\mathbf{x}}\rVert_{Z}\leq 2\eta\lVert\mathbf{g}\rVert_{Z^{-1}}.$
and
$\langle\mathbf{g},\widehat{\mathbf{x}}-\mathbf{x}^{*}\rangle\leq\frac{\eta}{2}\lVert\mathbf{g}\rVert_{Z^{-1}}^{2}+\frac{1}{2\eta}\lVert\widehat{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}-\frac{1}{2\eta}\lVert\widetilde{\mathbf{x}}-\mathbf{x}^{*}\rVert_{Z}^{2}.$

### v3

@@ G.1 Proof of Theorem 3
Our proof is similar to adversarial linear mixture MDPs with the unknown transition in Zhao et al. [2023]. We prove this lemma by reducing MNL mixture MDPs to a sequence of logistic bandits.
We use each three layers to construct a block. Note that the third layer of block $i$ is also the first layer of block $i+1$ and hence there are total $H/2$ blocks. In each block, both the first and third layers of this block only have one state, and the second layer has two states. Here we take block $i$ as an example. The first two layers of this block are associated with transition probability $\mathbb{P}_{i,1}$ and $\mathbb{P}_{i,2}$. Denote by $s_{i,1}$ the only state in the first layer of this block. In the second layer of the block $i$, we assume there exist two states $s_{i,2}^{*}$ and $s_{i,2}$. Let $s_{i,3}$ be the only state in the third layer of this block. Further, for any $a\in\mathcal{A}$, let $\phi(s_{i,1},a,s_{i,2})=0$. The transition probability is defined as follows:
$\mathbb{P}_{i,1}(s_{i,2}^{*}\mid s_{i,1},a)$ $=\frac{\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}{1+\exp(\phi(s_{i,2}^{*}\>|\>s_{i,1},a)^{\top}\theta_{i,1}^{*})}=\rho_{a},\quad\mathbb{P}_{i,1}(s_{i,2}\mid s_{i,1},a)=1-\rho_{a}.$
For the second layer, it satisfies $\forall s=s_{i,2}^{*},s_{i,2}$, and $a\in\mathcal{A}$, $\mathbb{P}_{i,2}(s_{i,3}\mid s,a)=1$. The reward satisfies $r_{k}(s_{i,1},a)=0$ for the first layer and $r_{k}(s_{i,2}^{*},a)=1$, $r_{k}(s_{i,2},a)=0$ for the second for all $a\in\mathcal{A}$.
Then, consider the logistic bandit problem where a learner selects action $x\in\mathbb{R}^{d}$ and receives a reward $r_{k}$ sampled from the Bernoulli distribution with mean $\mu(x^{\top}\theta^{*})=(1+\exp(x^{\top}\theta^{*}))^{-1}$. By this configuration, we can see that learning in each block of MDP can be regarded as learning a $d$-dimensional logistic bandit problem with $A$ arms, where the arm set is $\phi(s_{i,2}^{*}\>|\>s_{i,1},a)$ and the expected reward of each arm is $\rho_{a}$. Thus, learning this MNL mixture MDP equals to learning $H/2$ logistic bandit problems. This finishes the proof. ∎
Appendix H Supporting Lemmas
@@
$V_{t}=V+\sum_{s=1}^{t-1}X_{s}X_{s}^{\top},\quad S_{t}=\sum_{s=1}^{t-1}\eta_{s}X_{s}.$
Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$, for all $t\geq 1$,
$\left\|S_{t}\right\|_{V_{t}^{-1}}\leq R\sqrt{2\log\bigg{(}\frac{\operatorname{det}\left(V_{t}\right)^{1/2}\operatorname{det}(V)^{-1/2}}{\delta}\bigg{)}}.$
Lemma 13 (Périvier and Goyal [2022, Theorem 4]).
Let $\{\mathcal{F}_{t}\}_{t=0}^{\infty}$ be a filtration. Let $\{\delta_{t}\}_{t=1}^{\infty}$ be an $\mathbb{R}^{N}$-valued stochastic process such that $\delta_{t}$ is $\mathcal{F}_{t}$-measurable one-hot vector. Furthermore, assume $\mathbb{E}[\delta_{t}|\mathcal{F}_{t-1}]=p_{t}$ and define $\varepsilon_{t}=p_{t}-\delta_{t}$. Let $\{X_{t}\}_{t=1}^{\infty}$ be a sequence of $\mathbb{R}^{N\times d}$-valued stochastic process such that $X_{t}$ is $\mathcal{F}_{t-1}$-measurable and $\lVert X_{t,i}\rVert_{2}\leq 1,\forall i\in[N]$. Let $\{\lambda_{t}\}_{t=1}^{\infty}$ be a sequence of non-negative scalars. Define
$H_{t}=\sum_{i=1}^{t-1}\left(\sum_{j=1}^{N}p_{j}X_{i,j}X_{i,j}^{\top}-\sum_{j=1}^{N}\sum_{k=1}^{N}p_{j}p_{k}X_{i,j}X_{i,k}^{\top}\right)+\lambda_{t}I_{d},\quad S_{t}=\sum_{i=1}^{t-1}\sum_{j=1}^{N}\varepsilon_{i,j}X_{i,j}.$
Then, for any $\zeta\in(0,1)$, with probability at least $1-\zeta$, for all $t\geq 1$,
$\lVert S_{t}\rVert_{H_{t}^{-1}}\leq\frac{\sqrt{\lambda_{t}}}{4}+\frac{4}{\sqrt{\lambda_{t}}}\log\left(\frac{2^{d}\det\left(H_{t}\right)^{\frac{1}{2}}\lambda_{t}^{-\frac{d}{2}}}{\zeta}\right).$
@@
Suppose $x_{1},\ldots,x_{t}\in\mathbb{R}^{d}$ and for any $1\leq s\leq t$, $\lVert x_{s}\rVert_{2}\leq L$. Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$. Then, we have $\det(V_{t})\leq\left(\lambda+{tL^{2}}/{d}\right)^{d}$.
@@
Let $Z$ be a positive define matrix and $\mathcal{W}$ be a convex set, define $\mathbf{w}_{t+1}$ as the solution of
$\mathbf{w}_{t+1}=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{W}}\Big{\{}\langle\mathbf{g},\mathbf{w}\rangle+\frac{1}{2\eta}\lVert\mathbf{w}-\mathbf{w}_{t}\rVert_{Z}^{2}\Big{\}}.$
Then we have
$\lVert\mathbf{w}_{t+1}-\mathbf{w}_{t}\rVert_{Z}\leq 2\eta\lVert\mathbf{g}\rVert_{Z^{-1}}.$
Lemma 16 (Campolongo and Orabona [2020, Proposition 4.1]).
Define $\mathbf{w}_{t+1}$ as the solution of
$\mathbf{w}_{t+1}=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{V}}\big{\{}\eta\ell_{t}(\mathbf{w})+\mathcal{D}_{\psi}\left(\mathbf{w},\mathbf{w}_{t}\right)\big{\}},$
[…]
$\left\langle\eta_{t}\mathbf{g}_{t}^{\prime},\mathbf{w}_{t+1}-\mathbf{u}\right\rangle\leq\left\langle\nabla\psi\left(\mathbf{w}_{t}\right)-\nabla\psi\left(\mathbf{w}_{t+1}\right),\mathbf{w}_{t+1}-\mathbf{u}\right\rangle$
for any $\mathbf{u}\in\mathcal{W}$.